# EFC++: Elastic Feature Consolidation with Prototype Re-balancing for Cold Start Exemplar-free Incremental Learning

### v2

@@ VI Experimental Settings
VI-A Datasets
We evaluate EFC++ on five distinct datasets, with their statistics detailed in Appendix D-A. We categorize CIFAR-100 [65], Tiny-ImageNet [66], and ImageNet-Subset [67] as small-scale datasets, containing 100, 200, and 100 classes, respectively. For a more comprehensive large-scale scenario, we use ImageNet-1K [67], which includes 1000 diverse object classes with varying numbers of samples per class. Additionally, we utilize DN4IL, a well-crafted and balanced subset of the standard DomainNet dataset [32]. DN4IL consists of 600 classes, with 100 classes per domain, specifically clipart, infograph, painting, quickdraw, real, and sketch. This dataset is instrumental in comparing our approach to the state-of-the-art under large domain shifts.
VI-B Warm and Cold Start Scenarios
TABLE I: Class Distribution in Warm Start and Cold Start: $|\mathcal{C}_{0}|$ is the number of classes in the pre-training phase, while $|\mathcal{C}_{i>0}|$ represents classes introduced per incremental step.
@@ 10 Step 20 Step 10 Step 20 Step
We consider two scenarios for evaluating all approaches. The first, which we refer to as the Warm Start (WS) scenario which is commonly considered in the EFCIL literature [21, 22, 26, 23], involves pre-training the model on $50\%$ or $40\%$ of all available classes, denoted as $\mathcal{C}_{0}$. The remaining classes are then equally distributed among the subsequent incremental learning steps. This setting, which begins with a larger task favors continual learning techniques that prioritize stability in retained knowledge over the plasticity required to learn new classes.
The second scenario, which we call Cold Start (CS), is relatively unexplored in the the recent EFCIL literature. In this scenario, the model is randomly initialized before incremental learning, and all available classes are equally split among the incremental learning steps, resulting in no initial pre-training phase. Consequently, the number of classes per step is larger compared to Warm Start. See Table I for a detailed description of how the classes are distributed in Warm Start and Cold Start across all datasets.
We are primarily interested in the Cold Start scenario since it demands high backbone plasticity and poses a significant challenge for EFCIL. Mitigating forgetting becomes difficult without using exemplars when larger backbone shifts occur (in Section VIII-B we provide an empirical analysis of this phenomenon).
VI-C Metrics
We use as metrics the per-step incremental accuracy $A_{\text{step}}^{K}$ and average incremental accuracy $A_{\text{inc}}^{K}$:
$A_{\text{step}}^{K}=\frac{\sum_{i=s}^{K}|\mathcal{C}_{i}|a^{K}_{i}}{\sum_{i=s}^{K}|\mathcal{C}_{i}|},\,\,\,\,A_{\text{inc}}^{K}=\dfrac{1}{K}\sum_{i=s}^{K}A_{\text{step}}^{i},$ (21)
where $a^{K}_{i}$ represents the accuracy on task $i$ after training task $K$, and $s=1$ in the Cold Start scenario and $s=0$ in the Warm Start scenario. It should be noted that in the Warm Start (WS) scenario, the performance on the initial classes, indicated as step $s=0$, is commonly included in the evaluation metric [21, 23, 22, 26, 27], and we follow this same approach. Therefore, achieving higher stability on the large number of initial classes significantly impacts the final evaluation. In contrast, in the Cold Start (CS) scenario the classes are equally distributed among the incremental learning tasks and the evaluation metrics starting from $s=1$ give equal importance to each step.
To gain a more comprehensive understanding of impact of different regularization techniques employed by the tested methods, we use average forgetting $F^{K}$ to measure stability during training and cumulative average accuracy on the last task $\text{PL}^{K}$ to measure plasticity (i.e. the average of the diagonal of the matrix in Figure 4[68]. More specifically, the forgetting of a task is the difference between its maximum accuracy obtained in the past and its current accuracy:
@@ VI-C Metrics
$F^{K}=\dfrac{1}{K-1}\sum_{j=1}^{K-1}f_{j}^{K},\,\,\,\,\text{PL}^{K}=\dfrac{\sum_{i=s}^{K}|\mathcal{C}_{i}|a^{i}_{i}}{\sum_{i=s}^{K}|\mathcal{C}_{i}|}.$ (23)
VI-D Hyperparameter settings
We use the standard ResNet-18 [1] trained from scratch for all experiments. We train the first task of each state-of-the-art method using the same optimizer, number of epochs, and data augmentations. In particular, we train the first task using self-rotation as in [21, 23] to align the performance of all techniques.
For the incremental steps of EFC++ we used Adam a with a weight decay of 2e-4 and a fixed learning rate of 1e-4 for Tiny-ImageNet and CIFAR-100, while for ImageNet-Subset, ImageNet-1K, DN4IL we use a learning rate of 1e-5 for the backbone and 1e-4 for the heads. We fixed the total number of epochs to 100 and use a batch size of 64, except for ImageNet-1K where the batch size is fixed to 256. We set $\lambda_{\text{EFM}}=10$ and $\eta=0.1$ in Eq. 10 for all the experiments. In Section VIII-C, we discuss the theoretical basis for selecting these hyperparameters and analyze their impact on performance. For EFC++, during post-training we use SGD with a fixed learning rate 1e-3 to train the final classifier on both prototype and current task features for 50 epochs.
We ran all approaches using five random seeds and shuffling the classes in order to reduce the bias induced by the choice of class ordering [40, 14]. In Appendix D-B we provide the optimization settings for each state-of-the art method we evaluated.
VII Experimental Results
@@
VII-A Small-scale Class-IL Experiments
TABLE II: Small-scale Class-IL experiments. Fusion results are those reported in [23] as no code is provided to reproduce them. The result for ImageNet-Subset Cold Start for ABD and R-DFCIL are those reported in [30].
@@ VII-A Small-scale Class-IL Experiments
CIFAR-100 EWC $21.08\pm 1.09$ $13.53\pm 1.11$ $41.00\pm 1.11$ $31.79\pm 2.77$ $31.17\pm 2.94$ $17.37\pm 2.43$ $49.14\pm 1.28$ $31.02\pm 1.15$
LwF [^3] LwF results are not aligned with those in R-DFCIL, due to different architecture and implementation. We use the FACIL [14] implementation on ResNet-18, freezing old task classifier head while learning new task. $20.73\pm 1.47$ $11.78\pm 0.60$ $41.95\pm 1.30$ $28.93\pm 1.62$ $32.80\pm 3.08$ $17.44\pm 0.73$ $53.91\pm 1.67$ $38.39\pm 1.05$
@@
Tiny-ImageNet EWC $\phantom{0}6.73\pm 0.44$ $\phantom{0}5.96\pm 1.17$ $18.48\pm 0.60$ $13.74\pm 0.52$ $\phantom{0}8.00\pm 0.27$ $\phantom{0}5.16\pm 0.54$ $24.01\pm 0.51$ $15.70\pm 0.35$
LwF [^3] LwF results are not aligned with those in R-DFCIL, due to different architecture and implementation. We use the FACIL [14] implementation on ResNet-18, freezing old task classifier head while learning new task. $24.00\pm 1.44$ $\phantom{0}7.58\pm 0.37$ $43.15\pm 1.02$ $22.89\pm 0.64$ $26.09\pm 1.29$ $15.02\pm 0.67$ $45.14\pm 0.88$ $32.94\pm 0.54$
@@
Fusion $46.92$ $44.61$- -- -- -- -- --
FeTrIL $45.71\pm 0.39$ $44.63\pm 0.49$ $53.95\pm 0.42$ $52.96\pm 0.45$ $30.97\pm 0.90$ $25.70\pm 0.61$ $45.60\pm 1.67$ $39.54\pm 1.19$
SSRE $44.66\pm 0.45$ $44.68\pm 0.36$ $53.27\pm 0.43$ $52.94\pm 0.42$ $22.93\pm 0.95$ $17.34\pm 1.06$ $38.82\pm 1.99$ $30.62\pm 1.96$
@@
EFC 50.40 ±$\pm 0.25$ $48.68\pm 0.65$ $57.52\pm 0.43$ $56.52\pm 0.53$ $34.10\pm 0.77$ $\underline{28.69}\pm 0.40$ $47.95\pm 0.61$ $42.07\pm 0.96$
EFC++ 51.67 ±$\pm 0.31$ 50.41 ±$\pm 0.51$ 58.54 ±$\pm 0.51$ 57.79 ±$\pm 0.54$ 37.48 ±$\pm 0.52$ 32.56 ±$\pm 0.44$ 50.18 ±$\pm 1.11$ 45.58 ±$\pm 1.32$
ImageNet-Subset EWC $16.19\pm 2.48$ $10.66\pm 1.74$ $23.58\pm 2.01$ $18.05\pm 1.10$ $24.59\pm 4.13$ $12.78\pm 1.95$ $39.40\pm 3.05$ $26.95\pm 1.02$
LwF [^3] LwF results are not aligned with those in R-DFCIL, due to different architecture and implementation. We use the FACIL [14] implementation on ResNet-18, freezing old task classifier head while learning new task. $21.89\pm 0.52$ $13.24\pm 1.61$ $37.15\pm 2.47$ $25.96\pm 0.95$ $37.71\pm 2.53$ $18.64\pm 1.67$ $56.41\pm 1.03$ $40.23\pm 0.43$
@@
In Table II we compare EFC++ with baselines and state-of-the-art EFCIL approaches. Specifically, we consider EWC [10], LwF [53], PASS [21], Fusion [23], FeTrIL [26], SSRE [22], FeCAM [27], ABD [31], R-DFCIL [30], and EFC [29] on both Warm and Cold Start scenarios for varying numbers of incremental steps. EFC++ significantly outperforms the previous state-of-the-art in both metrics across all scenarios and datasets.
@@
Note that the incremental accuracy of SSRE and FeCAM in the Warm Start scenario differs from the values reported in the original paper, with SSRE higher and FeCAM lower. This discrepancy is attributable to the data augmentation policy we used and the self-rotation task we used initially to align all results with PASS and Fusion (both of which use self-rotation). This alignment is crucial when dealing with Warm Start scenarios, as performance on the large first task heavily biases the metrics (Eq. 21) as discussed in [69, 70]. In Appendix J-A and J-B we provide per-step performance plots in all the analyzed scenarios.
@@
Warm Start. The experimental results on Warm Start scenarios (Table II, left) reveal that EFC++ exhibits substantial improvements (ranging from 2 to 6 percentage points), over recent methods like PASS, Fusion, FeTrIL, and SSRE across all benchmarks. In this setting,  generative inversion-based methods (ABD, R-DFCIL) perform significantly worse than feature distillation approaches (PASS, Fusion, SSRE) and backbone freezing methods (FeTrIL, FeCAM). Compared to the newest and strongest competitor, FeCAM, EFC++ generally achieves better or comparable performance across all scenarios. Moreover, by comparing EFC++ and EFC, we see that the performance improvement is consistent regardless of the number of incremental learning steps. Specifically, for both CIFAR-100 and Tiny-ImageNet, in terms of per-step accuracy, EFC++ improves over EFC by approximately 1.3% and 1.7% for 10 and 20 steps, respectively; while for ImageNet-Subset EFC++ slightly improves the performance of EFC for both 10 and 20 incremental learning steps. This result is due that the drawback caused by the double backward pass required for computing $\mathcal{L}_{t}^{\text{PR-ACE}}$ in EFC, an effect reduced in the Warm Start scenario. This happens because the stronger backbone mitigates the overall drift, allowing EFC to perform almost as well as EFC$+$ in Warm Start (see Section VIII-B for drift comparison in WS and CS).
@@
When comparing EFC++ with FeCAM, we observe that EFC++ achieves comparable performance on CIFAR-100, while on Tiny-ImageNet and ImageNet-Subset, it improves performance by approximately 1.2% in the 10-step scenario and remains comparable in the 20-step scenario. This is expected, as FeCAM, which freezes the backbone after the first task, is specifically designed for WS and benefits from this constraint by preserving high first-task accuracy, which influences the final per-step accuracy. In contrast, our approach, which trains the feature extractor and enhances plasticity with the EFM regularizer, faces greater challenges in maintaining first-task performance, especially as the number of incremental steps increases. As a result, the per-step accuracy reflects this performance drop while remaining competitive with FeCAM (see Figure 6 for each task performance).
@@
Cold Start. In Cold Start scenarios, (Table II, right), methods relying on feature distillation, such as SSRE and PASS, or those that freeze the feature extractor, such as FeTrIL and FeCAM, experience a significant decrease in accuracy compared to EFC++. This drop in accuracy is due the fact that the first task does not provide a sufficiently strong starting point for the entire class-incremental process. Moreover, most of these approaches exhibit weaker performance even when compared to LwF, which does not use prototypes to balance the classifiers.
@@
EFC++ consistently and significantly improves the performance of EFC across all benchmarks and incremental learning steps. The key reason is that, in the Cold Start scenario – where the backbone is trained with a small number of classes – feature drift is significant (see Section VIII-B for the drift comparison in WS and CS). In this setting, EFC++ better controls drift and achieves approximately 3-4 percentage points higher accuracy. Additionally, the improvement of EFC++ over EFC can be attributed to the decoupling of backbone training and the balancing of classification heads. This ensures that the prototypes used in EFC++ do not become outdated during training, unlike in EFC.
@@
Figure 6 shows task-wise accuracy after learning the final task on CIFAR-100 in both WS and CS. Task 0 represents the large first task in WS, which is absent in CS (see Table I). EFC++ achieves the best accuracy on new tasks, highlighting the effectiveness of the EFM regularizer in enhancing plasticity. While its first-task accuracy is lower than FeCAM in both settings, it remains competitive with R-DFCIL in WS and surpasses it in CS, reinforcing its overall stability-plasticity balance. It should be noted that R-DFCIL struggles on new tasks in WS, despite its strong plasticity in CS. We attribute this to the sampling strategy of R-DFCIL for retrieving old task data from the generative inversion-based model, which can struggle when the number of classes in subsequent tasks is significantly lower than in the first one, as in Warm Start.
VII-B Large-scale Domain- and Class-IL Experiments
TABLE III: Large scale Class- and Domain-IL Experiments on ImageNet-1K (CS) and DN4IL [32] (CS) using $A_{\text{step}}^{K}$.

### v3

@@ 6 Experimental Settings
Datasets. We evaluate EFC++ on five distinct datasets, with their statistics detailed in Appendix D.1. We categorize CIFAR-100 CIFAR100 , Tiny-ImageNet TinyImageNet , and ImageNet-Subset ImageNet as small-scale datasets, containing 100, 200, and 100 classes, respectively. For a more comprehensive large-scale scenario, we use ImageNet-1K ImageNet , which includes 1000 diverse object classes with varying numbers of samples per class. Additionally, we utilize DN4IL, a well-crafted and balanced subset of the standard DomainNet dataset gowda2023dual . DN4IL consists of 600 classes, with 100 classes per domain, specifically clipart, infograph, painting, quickdraw, real, and sketch. This dataset is instrumental in comparing our approach to the state-of-the-art under large domain shifts.
Table 1: Class Distribution in Warm Start and Cold Start: $|\mathcal{C}_{0}|$ is the number of classes in the pre-training phase, while $|\mathcal{C}_{i>0}|$ represents classes introduced per incremental step.
@@ 10 Step 20 Step 10 Step 20 Step
Warm and Cold Start. We consider two scenarios for evaluating all approaches. The first, which we refer to as the Warm Start (WS) scenario which is commonly considered in the EFCIL literature pass ; ssre ; fetril ; evanescent , involves pre-training the model on $50\%$ or $40\%$ of all available classes, denoted as $\mathcal{C}_{0}$. The remaining classes are then equally distributed among the subsequent incremental learning steps. This setting, which begins with a larger task favors continual learning techniques that prioritize stability in retained knowledge over the plasticity required to learn new classes.
The second scenario, which we call Cold Start (CS), is relatively unexplored in the the recent EFCIL literature. In this scenario, the model is randomly initialized before incremental learning, and all available classes are equally split among the incremental learning steps, resulting in no initial pre-training phase. Consequently, the number of classes per step is larger compared to Warm Start. See Table 1 for a detailed description of how the classes are distributed in Warm Start and Cold Start across all datasets.
We are primarily interested in the Cold Start scenario since it demands high backbone plasticity and poses a significant challenge for EFCIL. Mitigating forgetting becomes difficult without using exemplars when larger backbone shifts occur (in Section 8.2 we provide an empirical analysis of this phenomenon).
Metrics. We use as metrics the per-step incremental accuracy $A_{\text{step}}^{K}$ and average incremental accuracy $A_{\text{inc}}^{K}$:
$A_{\text{step}}^{K}=\frac{\sum_{i=s}^{K}|\mathcal{C}_{i}|a^{K}_{i}}{\sum_{i=s}^{K}|\mathcal{C}_{i}|},\,\,A_{\text{inc}}^{K}=\dfrac{1}{K-s+1}\sum_{i=s}^{K}A_{\text{step}}^{i},$ (21)
where $a^{K}_{i}$ represents the accuracy on task $i$ after training task $K$, and $s=1$ in the Cold Start scenario and $s=0$ in the Warm Start scenario. It should be noted that in the Warm Start (WS) scenario, the performance on the initial classes, indicated as step $s=0$, is commonly included in the evaluation metric pass ; evanescent ; ssre ; fetril ; goswami2023fecam , and we follow this same approach. Therefore, achieving higher stability on the large number of initial classes significantly impacts the final evaluation. In contrast, in the Cold Start (CS) scenario the classes are equally distributed among the incremental learning tasks and the evaluation metrics starting from $s=1$ give equal importance to each step.
To gain a more comprehensive understanding of impact of different regularization techniques employed by the tested methods, we use average forgetting $F^{K}$ to measure stability during training and cumulative average accuracy on the last task $\text{PL}^{K}$ to measure plasticity (i.e. the average of the diagonal of the matrix in Figure 4lange2023continual . More specifically, the forgetting of a task is the difference between its maximum accuracy obtained in the past and its current accuracy:
@@
$F^{K}=\dfrac{1}{K-s}\sum_{j=s}^{K-1}f_{j}^{K},\,\,\,\,\text{PL}^{K}=\frac{1}{K-s+1}\sum_{i=s}^{K}a^{i}_{i}.$ (23)
Hyperparameter settings. We use the standard ResNet-18 resnet trained from scratch for all experiments. We train the first task of each state-of-the-art method using the same optimizer, number of epochs, and data augmentations. In particular, we train the first task using self-rotation as in pass ; evanescent to align the performance of all techniques.
For the incremental steps of EFC++ we used Adam with a weight decay of 2e-4 and a fixed learning rate of 1e-4 for Tiny-ImageNet and CIFAR-100, while for ImageNet-Subset, ImageNet-1K, DN4IL we use a learning rate of 1e-5 for the backbone and 1e-4 for the heads. We fixed the total number of epochs to 100 and use a batch size of 64, except for ImageNet-1K where the batch size is fixed to 256. We set $\lambda_{\text{EFM}}=10$ and $\eta=0.1$ in Eq. 10 for all the experiments. In Section 8.3, we discuss the theoretical basis for selecting these hyperparameters and analyze their impact on performance. For EFC++, during post-training we use SGD with a fixed learning rate 1e-3 to train the final classifier on both prototype and current task features for 50 epochs.
We ran all approaches using five random seeds and shuffling the classes in order to reduce the bias induced by the choice of class ordering icarl ; facil . In Appendix D.2 we provide the optimization settings for each state-of-the art method we evaluated.
7 Experimental Results
@@
7.1 Small-scale Class-IL Experiments
Table 2: Small-scale Class-IL experiments. Fusion results are those reported in evanescent as no code is provided to reproduce them. The result for ImageNet-Subset Cold Start for ABD and R-DFCIL are those reported in gao2022r .
@@ 7.1 Small-scale Class-IL Experiments
CIFAR-100
EWC $21.08\pm 1.09$ $13.53\pm 1.11$ $41.00\pm 1.11$ $31.79\pm 2.77$ $31.17\pm 2.94$ $17.37\pm 2.43$ $49.14\pm 1.28$ $31.02\pm 1.15$
LwF [^3] LwF results are not aligned with those in R-DFCIL, due to different architecture and implementation. We use the FACIL facil implementation on ResNet-18, freezing old task classifier head while learning new task. $20.73\pm 1.47$ $11.78\pm 0.60$ $41.95\pm 1.30$ $28.93\pm 1.62$ $32.80\pm 3.08$ $17.44\pm 0.73$ $53.91\pm 1.67$ $38.39\pm 1.05$
@@
Tiny-ImageNet
EWC $\phantom{0}6.73\pm 0.44$ $\phantom{0}5.96\pm 1.17$ $18.48\pm 0.60$ $13.74\pm 0.52$ $\phantom{0}8.00\pm 0.27$ $\phantom{0}5.16\pm 0.54$ $24.01\pm 0.51$ $15.70\pm 0.35$
LwF [^3] LwF results are not aligned with those in R-DFCIL, due to different architecture and implementation. We use the FACIL facil implementation on ResNet-18, freezing old task classifier head while learning new task. $24.00\pm 1.44$ $\phantom{0}7.58\pm 0.37$ $43.15\pm 1.02$ $22.89\pm 0.64$ $26.09\pm 1.29$ $15.02\pm 0.67$ $45.14\pm 0.88$ $32.94\pm 0.54$
@@
Fusion $46.92$ $44.61$ - - - - - -
FeTrIL $5.71\pm 0.39 44$$4.63\pm 0.49$ $53.95\pm 0.42$ $52.96\pm 0.45$ $30.97\pm 0.90$ $25.70\pm 0.61$ $45.60\pm 1.67$ $39.54\pm 1.19$
SSRE $4.66\pm 0.45 44$$4.68\pm 0.36$ $53.27\pm 0.43$ $52.94\pm 0.42$ $22.93\pm 0.95$ $17.34\pm 1.06$ $38.82\pm 1.99$ $30.62\pm 1.96$
@@
EFC 50.40 $\pm 0.25$ $48.68\pm 0.65$ $57.52\pm 0.43$ $56.52\pm 0.53$ $34.10\pm 0.77$ $\underline{28.69}\pm 0.40$ $47.95\pm 0.61$ $42.07\pm 0.96$
EFC++ 51.67 $\pm 0.31$ 50.41 $\pm 0.51$ 58.54 $\pm 0.51$ 57.79 $\pm 0.54$ 37.48 $\pm 0.52$ 32.56 $\pm 0.44$ 50.18 $\pm 1.11$ 45.58 $\pm 1.32$
ImageNet-Subset
EWC $16.19\pm 2.48$ $10.66\pm 1.74$ $23.58\pm 2.01$ $18.05\pm 1.10$ $24.59\pm 4.13$ $12.78\pm 1.95$ $39.40\pm 3.05$ $26.95\pm 1.02$
LwF [^3] LwF results are not aligned with those in R-DFCIL, due to different architecture and implementation. We use the FACIL facil implementation on ResNet-18, freezing old task classifier head while learning new task. $21.89\pm 0.52$ $13.24\pm 1.61$ $37.15\pm 2.47$ $25.96\pm 0.95$ $37.71\pm 2.53$ $18.64\pm 1.67$ $56.41\pm 1.03$ $40.23\pm 0.43$
@@
In Table 2 we compare EFC++ with baselines and state-of-the-art EFCIL approaches. Specifically, we consider EWC ewc , LwF lwf , PASS pass , Fusion evanescent , FeTrIL fetril , SSRE ssre , FeCAM goswami2023fecam , ABD smith2021always , R-DFCIL gao2022r , and EFC magistri2024elastic on both Warm and Cold Start scenarios for varying numbers of incremental steps. EFC++ significantly outperforms the previous state-of-the-art in both metrics across all scenarios and datasets.
@@
Note that the incremental accuracy of SSRE and FeCAM in the Warm Start scenario differs from the values reported in the original paper, with SSRE higher and FeCAM lower. This discrepancy is attributable to the data augmentation policy we used and the self-rotation task we used initially to align all results with PASS and Fusion (both of which use self-rotation). This alignment is crucial when dealing with Warm Start scenarios, as performance on the large first task heavily biases the metrics (Eq. 21) as discussed in pycl ; endtoend . In Appendix J.1 and J.2 we provide per-step performance plots in all the analyzed scenarios.
@@
Warm Start. The experimental results on Warm Start scenarios (Table 2, left) reveal that EFC++ exhibits substantial improvements (ranging from 2 to 6 percentage points), over recent methods like PASS, Fusion, FeTrIL, and SSRE across all benchmarks. In this setting, generative inversion-based methods (ABD, R-DFCIL) perform significantly worse than feature distillation approaches (PASS, Fusion, SSRE) and backbone freezing methods (FeTrIL, FeCAM). Compared to the newest and strongest competitor, FeCAM, EFC++ generally achieves better or comparable performance across all scenarios. Moreover, by comparing EFC++ and EFC, we see that the performance improvement is consistent regardless of the number of incremental learning steps. Specifically, for both CIFAR-100 and Tiny-ImageNet, in terms of per-step accuracy, EFC++ improves over EFC by approximately 1.3% and 1.7% for 10 and 20 steps, respectively; while for ImageNet-Subset EFC++ slightly improves the performance of EFC for both 10 and 20 incremental learning steps. This result is due that the drawback caused by the double backward pass required for computing $\mathcal{L}_{t}^{\text{PR-ACE}}$ in EFC, an effect reduced in the Warm Start scenario. This happens because the stronger backbone mitigates the overall drift, allowing EFC to perform almost as well as EFC++ in Warm Start (see Section 8.2 for drift comparison in WS and CS).
@@
When comparing EFC++ with FeCAM, we note that EFC++ achieves comparable performance on CIFAR-100, while on Tiny-ImageNet and ImageNet-Subset, it improves performance by approximately 1.2% in the 10-step scenario and remains comparable in the 20-step scenario. This is expected, as FeCAM, which freezes the backbone after the first task, is specifically designed for WS and benefits from this constraint by preserving high first-task accuracy, which influences the final per-step accuracy. In contrast, our approach, which trains the feature extractor and enhances plasticity with the EFM regularizer, faces greater challenges in maintaining first-task performance, especially as the number of incremental steps increases. As a result, the per-step accuracy reflects this performance drop while remaining competitive with FeCAM (see Figure 6 for each task performance).
@@
Cold Start. In Cold Start scenarios, (Table 2, right), methods relying on feature distillation, such as SSRE and PASS, or those that freeze the feature extractor, such as FeTrIL and FeCAM, experience a significant decrease in accuracy compared to EFC++. This drop in accuracy is due the fact that the first task does not provide a sufficiently strong starting point for the entire class-incremental process. Moreover, most of these approaches exhibit weaker performance even when compared to LwF, which does not use prototypes to balance the classifiers.
@@
EFC++ consistently and significantly improves the performance of EFC across all benchmarks and incremental learning steps. The key reason is that, in the Cold Start scenario – where the backbone is trained with a small number of classes – feature drift is significant (see Section 8.2 for the drift comparison in WS and CS). In this setting, EFC++ better controls drift and achieves approximately 3-4 percentage points higher accuracy. Additionally, the improvement of EFC++ over EFC can be attributed to the decoupling of backbone training and the balancing of classification heads. This ensures that the prototypes used in EFC++ do not become outdated during training, unlike in EFC.
@@
Figure 6 shows task-wise accuracy after learning the final task on CIFAR-100 in both WS and CS. Task 0 represents the large first task in WS, which is absent in CS (see Table 1). EFC++ achieves the best accuracy on new tasks, highlighting the effectiveness of the EFM regularizer in enhancing plasticity. While its first-task accuracy is lower than FeCAM in both settings, it remains competitive with R-DFCIL in WS and surpasses it in CS, reinforcing its overall stability-plasticity balance. It should be noted that R-DFCIL struggles on new tasks in WS, despite its strong plasticity in CS. We attribute this to the sampling strategy of R-DFCIL for retrieving old task data from the generative inversion-based model, which can struggle when the number of classes in subsequent tasks is significantly lower than in the first one, as in Warm Start.
7.2 Large-scale Domain- and Class-IL Experiments
Table 3: Large scale Class- and Domain-IL Experiments on ImageNet-1K (CS) and DN4IL gowda2023dual (CS) using $A_{\text{step}}^{K}$.